40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 9449     Entry time: Fri Dec 6 21:38:27 2013
Author: Koji 
Type: Update 
Category: LSC 
Subject: CDS related activities for LSC 

I worked on the CDS related stuffs for LSC yesterday and today.


1. Slow machines:

I checked the database files for c1iscaux and c1iscaux2 (slow machines). They are mainly
used for the control of LSC whitening filters. The channel names were totally random as we
reconfigured the RF PDs while the channel names had been unchanged.

- Now the database was modified so that the PD name and the channels are related.
- saverestore.req and autoBurt.req were also changed accordingly.

- PD interface channels are completely random. Don't use them.
- I found the whitening of DCPDs are not effective.

- We need to clean up /cvs/cds/caltech/target directory. The autoBurt requests in the old targets
are making unnecessary burt files.

2. LSC screens

- The channel names on the LSC OVERVIEW screen was modified. (Attachment 1)
- A new LSC Whitening screen was made. (Attachment 2)

3. LSC screen generator

To touch the main LSC screen is very tough. The screen was split in to several sub screens
and combined with a command.

/opt/rtcds/caltech/c1/medm/c1lsc/master/generateLSCscreen/generateLSCscreen.py

This command combines the multiple adl files into a single file with x&y offsets.
This way, you can work with the each section of the screen.
Also, moving the blocks are just easy.

4. LSC Code Bug?

During the screen making, I found that a couple of the whitening switches are not
working properly.
e.g. When AS165 (either I or Q) FM1 is activated throught the whitening trigger,
the MSB bit (bit15) of the binary I/O (C1:LSC-BIO_0_0) does not .

SImilarly ASDC FM1 does not toggle bit15 of C1:LSC-BIO_0_1.

The other channels seems OK.

At first, I thought this is a bug of "Bit2Word" block. But an individual test of the block showed that
the block is not guilty. So why is only Bit15 malfunctioning???

 

Attachment 1: LSC1.png  38 kB  | Hide | Hide all
LSC1.png
Attachment 2: LSC2.png  31 kB  | Hide | Hide all
LSC2.png
ELOG V3.1.3-